Oxygen Optimization is considering the caffeine project, which would involve selling caffeinated oxygen for 1 year. The firm expects sales of caffeinated oxygen to be 68,000 dollars and associated costs from providing caffeinated oxygen (such as tanks, filters, etc.) to be 33,000 dollars. The firm believes that sales of regular oxygen, which is currently sold by the firm, would be 37,000 dollars less with the addition of caffeinated oxygen, and that costs associated with regular oxygen to be 25,000 less with the addition of the caffeinated oxygen. Finally, Oxygen Optimization believes that the introduction of caffeinated oxygen would increase traffic to its facilities, which would increase expected sales of other products (such as masks) by 17,500 dollars more than it would be without the addition of caffeinated oxygen, and increase costs by 12,000 more than it would be without the addition of caffeinated oxygen. What is the operating cash flow (OCF) for year 1 that Oxygen Optimization should use to analyze the caffeine project? The tax rate is 30 percent and the cost of capital is 12.09 percent. Relevant depreciation is expected to be 9,000 dollars.
Answer #1

Sales1=68000

sales 68000 PV Of Sales=68000/(1.1209)=60665.5366
Decrease in sales of regularoxygen -37000=Pv=37000/(1.1209)=-33009.189
Increase in sales of other products 17500=Pv=17500/(1.1209)=15612.4543
Total benefit from sales =43268.8019

Operating cost

33000 PV=33000/(1.1209)=29440.6281
Decrease in operating Cost 25000 PV=-22303.5061
Increase in Operating cost 12000 PV=10705.6829
Total operating Cost 17842.8049

EBIT=Total sales-Operating cost=43268.8019-17842.8049=25425.997.

EBIT-Deprciation=25425.997-9000=16425.997

EBT-tax=16425.997-(4927.7991)=11498.1979.

Operating cash flow=EAT+Depreciation(NON CASH EXPENSES)=11498.1979+9000=20498.1979 dollars